# Exploring Unseen Environments with Robots using Large Language and Vision Models through a Procedurally Generated 3D Scene Representation

### v1

@@ Abstract
Recent advancements in Generative Artificial Intelligence, particularly in the realm of Large Language Models (LLMs) and Large Vision Language Models (LVLMs), have enabled the prospect of leveraging cognitive planners within robotic systems. This work focuses on solving the object goal navigation problem by mimicking human cognition to attend, perceive and store task specific information and generate plans with the same. We introduce a comprehensive framework capable of exploring an unfamiliar environment in search of an object by leveraging the capabilities of Large Language Models(LLMs) and Large Vision Language Models (LVLMs) in understanding the underlying semantics of our world. A challenging task in using LLMs to generate high level sub-goals is to efficiently represent the environment around the robot. We propose to use a 3D scene modular representation, with semantically rich descriptions of the object, to provide the LLM with task relevant information. But providing the LLM with a mass of contextual information (rich 3D scene semantic representation), can lead to redundant and inefficient plans. We propose to use an LLM based pruner that leverages the capabilities of in-context learning to prune out irrelevant goal specific information.
@@
I INTRODUCTION
Navigation in an unfamiliar environment to search for an object described in natural language is one of the most challenging problem in robotics. Even though these kind of task comes as second nature for us humans, the underlying process is really complex as it involves cognitive processing, using long term memory and experiences, and integrating the current sensory information with these processes. Achieving this with robots involve designing a language conditioned high-level planner that understands the semantic priors of this world and is able to devise intelligent sequential plans to efficiently explore the scene in search of the object. This is the core of an object goal navigation task. The agent, similar to a human being, must also have a cognitive high level-planner that can understand the semantic cues of our world, a low-level planner that can leverage on long term memory and experiences to act on the environment and an episodic memory to save critical information related to the current task.
Large Language Models (LLMs) and Large Vision Language Models (LVLMs) have shown the capability to comprehend the semantic priors of the world and reason about them from this understanding. This makes them an ideal candidate for a high level planner, that can take an action by reasoning about the scene. To plan efficiently in an environment by taking into account the underlying details of the scene, the agent should be provided with an efficient representation of the scene which is rich in semantic information but yet not overwhelmed with a mass of indistinguishable stimuli. Human beings, when exploring an unknown environment to find an object without a comprehensive map of the environment, follows a goal oriented approach in storing data. This process involves a combination of Perception which process the incoming information, Attention which directs our focus to specific cues in the scene and memory which selectively encodes information that is perceived as important for achieving a specific goal. This aids in efficient use of cognitive resources and effective retrieval of this information when needed. This goal oriented approach enhances the chances of us finding the target object. We mimic this goal oriented approach with robots by using a combination of an Large language model for attending to relevant aspects, an large vision language model and an open-vocabulary image segmentation module for perceiving and processing the incoming information, and a 3D scene modular representation and a short term memory module for saving the processed information to memory.
@@ I INTRODUCTION
The main contributions of our work involve:
  • A framework inspired from human cognition to attend, perceive and store information, to find an object in an unknown environment, that leverages the reasoning capability of Large Language Models (LLMs) and Large Vision Language models (LVLMs), to deduce efficient plans by procedurally building a 3D scene modular representation.
  • In-context learning based approach with LLMs to find the pertinent objects in the scene, to build a goal oriented, streamlined and scalable 3D scene representation on the fly, which aids in effective reasoning and efficient exploration in an unfamiliar environment.
  • An inference strategy that uses an LVLM to identify whether a detected segment is the target object or not, on all the frames in the short term memory, in which this detected segment was visible.
II Related works
II-A LLM as a planner
@@
Large language models (LLMs) with billions of parameters, trained on massive scale datasets have shown impressive generative capabilities with a generalized semantic understanding of the world. Many works like [1, 2, 3, 4, 5, 6] focused on grounding LLMs for high-level task planning with natural language task representation. [5] worked on grounding high-level tasks expressed in natural language and decomposing them into low level plans and admissible actions. [6] focused on grounding an LLM that provides high-level procedures for task completion, with value functions associated with these tasks. Many approaches like [7] and [8] used LLMs to follow natural language navigation instructions.
II-B Open Vocabulary Image Segmentation
@@
Open vocabulary image segmentation refers to the recognition and delineation of an open category of objects in a frame. The flexibility of such a system to identify and segment objects forms a critical part in building an efficient representation of the scene. With the advent of powerful object detection and segmentation models like [9, 10, 11, 12, 13, 14, 15], the agent can detect and segment the objects in the scene. Many open vocabulary segmentation modules like [12, 14, 9] require the labels of the objects to perform instance segmentation. Image tagging models like [16] can recognize the objects in a scene and return their labels, which can be used as text prompt to these models.
II-C Representating the environment for task planning
@@
Prior works like [17, 18, 8, 19, 20, 21] focused on building efficient representations of the environment to facilitate downstream task planning. The representations generated can be used for understanding the semantics of the scene, thus allowing the agent to query the environment for an object using natural language [17, 18] and generate task level plans [22, 18] or low-level actions [21]. Even though many scene representations like [19, 21] have created a scalable and efficient representation that can be used for navigation, it lacks contextual information about the scene, which is invaluable when exploring a scene. Scene representations like [17, 18] have a really dense representation, which is often not required when exploring an unfamiliar environment in search of a target object $G_{o}$. Our present research is inspired by the human way of cognition with an affordance based memory that attends and stores only those information which are relevant to the task at hand. We leverage on the ability of In-context learning in LLMs to prune out irrelevant information and then using an LVLM to reason about the relevant pruned list of objects, thus creating a sparse as well as information rich representation of the environment. The scene representation near the target object $G_{o}$ is made denser so as to facilitate further downstream task planning like manipulation of the object.
II-D Object goal navigation
@@
Object goal navigation task involves finding a target object $G_{o}$ in an unfamiliar environment. Efficient exploration through the environment requires deep understanding of semantic priors of the world. Prior works like [19, 23, 24, 25, 26, 27] tried to learn these semantic cues from egocentric RGB and depth images [23, 24], semantic map [19] and even from YouTube videos [25]. These learned semantic cues might not generalize well to new unseen environments. [23] focused on using a modular transfer learning model to generalize the learned policy for a particular task to multiple tasks and [28] used meta learning to generalize to unfamiliar environments. LLMs on the other hand exhibit extraordinary contextual awareness and ability to understand the semantic cues of our world. We leverage on this capability to efficiently explore unseen environments for the target object $G_{o}$.
@@
Recent works like [29, 30, 31] have leveraged the capability of LLMs to understand the semantic priors for object goal navigation tasks. [29] leveraged the planning capabilities of LLMs to devise a sequential plan that includes point goals and target state information for executing multiple object goal navigation tasks in parallel. These point goals generated by an LLM is fulfilled through a low level(execution level) controller. [31] focused on using the semantic predictions from LLMs as a heuristic for a frontier based exploration strategy to find the target object $G_{o}$. [30] focused on solving language-driven zero-shot object goal navigation problem [32] by using an LLM to navigate to the target object $G_{o}$, given a natural language description of it.
@@
III METHOD
Figure 2: architecture of the proposed pipeline.
Our work aims to design a planner for object goal navigation task in an unfamiliar environment by building a semantic 3D scene representation of the environment and utilizing an LLM to generate plans on top of this representation. The goal of the agent is to find a target object $G_{o}$ in the environment and to create a rich representation of the scene around the object for downstream planning. We leverage the reasoning capability of Large Language Models(LLMs) to understand the semantic priors of the environment and thus to guide the robot to those regions which have a higher chance of finding the target object $G_{o}$. The agent constructs a semantic 3D scene modular representation which is a sparse version of [17], pruned with LLMs using the in-context learning paradigm [33, 34]. The immediate relevant environment of the target object $G_{o}$ will have a dense and rich representation while the rest will have a sparse representation. The overview of our approach is given in figure 2.
@@ III METHOD
The agent receives RGB image, depth image, base and camera pose as inputs in every timestep. An open vocabulary semantic segmentation module identifies objects in the RGB image and generates mask for the same. LLM acting as a pruner identifies the most important objects in the scene in regards to their utility in understanding the semantic priors of the environment and prunes the remaining detections. These detected segments are then converted into a 3D scene representation, where each object is signified by a node. An object node stores information regarding the position of the object, pointcloud, frame in which this object was detected, mask and label of this object in that frame, and a semantic description of the object in the scene. LLM, grounded for an object goal navigation task, uses the 3D scene modular representation to decide whether to continue exploring the environment or to move closer to an already detected object in the scene that has a high probability of finding the target object $G_{o}$. The agent triggers its short term memory module and stores information for every frame if it decides to explore closer to an object. Upon reaching the object, the agent utilize the detections from the short term memory, to construct hypotheses regarding the target object.
III-A Open vocabulary image segmentation
@@
The open vocabulary segmentation module segments the RGB frame $I_{t}$ at the current timestep $t$, given the natural language description of the object. Given an RGB image $I_{t}$ of the scene, we use a combination of Recognize anything model(RAM) [16], Grounding Dino and FastSAM [14] for open vocabulary semantic segmentation. RAM is a strong foundational model used for image tagging, it is tasked with finding the object tags $c_{t}^{i}$ in the scene. The set of object tags $C_{t}=(c_{t}^{1},c_{t}^{2},...,c_{t}^{m})$ is then passed as a text prompt $T_{t}$ $(=C_{t})$ to Grounding-Dino [15], which detects the presence of classes $C_{t}$ in the scene. Grounding-Dino returns the class labels $(l_{t}^{1},l_{t}^{2},...,l_{t}^{m})$ of objects in the scene, confidence of the predictions $(p_{t}^{1},p_{t}^{2},...,p_{t}^{m})$ and their bounding boxes $(b_{t}^{1},b_{t}^{2},...,b_{t}^{m})$. The detected classes in the scene $(l_{t}^{1},l_{t}^{2},...,l_{t}^{m})$ are then passed as box prompt $(b_{t}^{1},b_{t}^{2},...,b_{t}^{m})$ through FastSAM [13] to get their corresponding masks $(m_{t}^{1},m_{t}^{2},...,m_{t}^{m})$.
III-B LLM as a pruner
@@
We leverage the In-context learning abilities of an LLM to prune the detected class labels $(l_{t}^{1},l_{t}^{2},...,l_{t}^{m})$ generated by the open vocabulary segmentation module for the timestep $t$. In-context learning is a paradigm that empowers pre-trained LLMs to ground to a particular task without fine-tuning the model, by providing the LLM with task level demonstrations along with the prompt. The language model here is provided with task-specific examples containing a pair of input $P_{i}$ and output $P_{o}$ object sets. The input set $P_{i}$ contains class labels $(l^{1},l^{2},...,l^{m})$ and the output set $P_{o}(\subseteq P_{i})$, contains those objects that have a higher utility when it comes to understanding the semantic priors of the scene. The text prompt provided to the LLM to achieve this behaviour is given in fig 3.
Figure 3: prompt to LLM for pruning the detections.
III-C 3D scene modular representation
@@
The pruned semantic masks $(m_{t}^{1},m_{t}^{2},...,m_{t}^{m})$ and labels $(l_{t}^{1},l_{t}^{2},...,l_{t}^{m})$ from the open vocabulary segmentation module, along with the depth image $I_{t}^{D}$ and camera pose $P_{t}$ in the current time step $t$ are used to generate 3D object nodes $(N_{c1},N_{c2},...,N_{ck})$ for each object $O_{i}$ in that frame $(I_{t},I_{t}^{D})$ using a modified concept graphs pipeline [17]. In concept graphs, the 3D object nodes $N_{ci}$ in the scene representation are fused from multiple views considering their spatial overlap, using an object association strategy that compares the semantic and visual similarity of these objects. Merging objects from different frames often creates conflicts in class labels, as the object segment can be associated with different class labels in different frames. We propose to use a Large Vision Language Model (LVLM) to resolve this conflict. The LVLM is tasked to find the class label that best suits the object detected, by giving it a cropped image of the detected object and the conflicting class labels.
@@
III-D LLM as a planner
LLM, grounded in the current task using task level description of the scene is used to generate a high level plan for the agent. Given the current objects in the environment (objects in the constructed 3D scene representation ($N_{c1},N_{c2},...,N_{ck}$)) and the generated descriptions for these object nodes, the LLM is tasked to identify whether to continue exploring the environment randomly and build up the 3D scene representation, or to explore the regions closer to an object node $N_{cj}$. If the LLM decides to explore the scene more, then $action$ is set to $<explore\textunderscore scene>$ and if it believes that there is a high chance of finding the target object $G_{o}$ near any of the detected objects, then $action$ is set to $<explore\textunderscore obj>$. This ability of an LLM to extract semantic relationships between the target object $G_{o}$ and objects in the scene ($N_{c1},N_{c2},...,N_{ck}$) helps the agent to explore efficiently in the environment. The text prompt provided to the LLM to achieve this behaviour is given in fig 4.
Figure 4: prompt to LLM for planning.
III-E Short term memory(STM) and reasoning

### v2

# Cognitive Planning for Object Goal Navigation using Generative AI Models

@@ Abstract
Recent advancements in Generative AI, particularly in Large Language Models (LLMs) and Large Vision-Language Models (LVLMs), offer new possibilities for integrating cognitive planning into robotic systems. In this work, we present a novel framework for solving the object goal navigation problem that generates efficient exploration strategies. Our approach enables a robot to navigate unfamiliar environments by leveraging LLMs and LVLMs to understand the semantic structure of the scene. To address the challenge of representing complex environments without overwhelming the system, we propose a 3D modular scene representation, enriched with semantic descriptions. This representation is dynamically pruned using an LLM-based mechanism, which filters irrelevant information and focuses on task-specific data. By combining these elements, our system generates high-level sub-goals that guide the robot’s exploration toward the target object. We validate our approach in simulated environments, demonstrating its ability to enhance object search efficiency while maintaining scalability in complex settings. Video Demonstration : https://youtu.be/pvr1uaObL9M
Keywords: Cognitive planning, Generative AI, Object goal navigation
@@
1 Introduction
Navigation in an unfamiliar environment to search for an object described in natural language is one of the most challenging problem in robotics [1, 2]. Even though these kind of task comes as second nature for us humans, the underlying process is really complex as it involves cognitive processing, using long term memory and experiences, and integrating the current sensory information with these processes. Achieving this with robots involve designing a language conditioned high-level planner [3] that understands the semantic priors of this world and is able to devise intelligent sequential plans to efficiently explore the scene in search of the object. This is the core of an object goal navigation task. The agent, similar to a human being, must also have a cognitive high level-planner that can understand the semantic cues of our world, a low-level planner that can leverage on long term memory and experiences to act on the environment and an episodic memory to save critical information related to the current task.
Large Language Models (LLMs) and Large Vision Language Models (LVLMs) have shown the capability to comprehend the semantic priors of the world and reason about them from this understanding. This makes them an ideal candidate for a high level planner, that can take an action by reasoning about the scene. To plan efficiently in an environment by taking into account the underlying details of the scene, the agent should be provided with an efficient representation of the scene which is rich in semantic information but yet not overwhelmed with a mass of indistinguishable stimuli. Human beings, when exploring an unknown environment to find an object without a comprehensive map of the environment, follows a goal oriented approach in storing data [4]. This process involves a combination of Perception which process the incoming information, Attention which directs our focus to specific cues in the scene and memory which selectively encodes information that is perceived as important for achieving a specific goal. This aids in efficient use of cognitive resources and effective retrieval of this information when needed. This goal oriented approach enhances the chances of us finding the target object. We mimic this goal oriented approach with robots by using a combination of an Large language model for attending to relevant aspects, an large vision language model and an open-vocabulary image segmentation module for perceiving and processing the incoming information, and a 3D scene modular representation and a short term memory module for saving the processed information to memory.
@@ 1 Introduction
2 Related works
2.1 LLM as a planner
@@
Large language models (LLMs) with billions of parameters, trained on massive scale datasets have shown impressive generative capabilities with a generalized semantic understanding of the world. Many works like Huang et al. [5], Ding et al. [6], Song et al. [7], Lin et al. [8], Huang et al. [9], Ahn et al. [10] focused on grounding LLMs for high-level task planning with natural language task representation. Huang et al. [9] worked on grounding high-level tasks expressed in natural language and decomposing them into low level plans and admissible actions. Ahn et al. [10] focused on grounding an LLM that provides high-level procedures for task completion, with value functions associated with these tasks.
2.2 Open Vocabulary Image Segmentation
@@
Open vocabulary image segmentation refers to the recognition and delineation of an open category of objects in a frame. The flexibility of such a system to identify and segment objects forms a critical part in building an efficient representation of the scene. With the advent of powerful object detection and segmentation models like Kirillov et al. [11], Wang et al. [12], He et al. [13], Zhou et al. [14], Zhao et al. [15], Ren et al. [16], Liu et al. [17], the agent can detect and segment the objects in the scene. Many open vocabulary segmentation modules like Zhou et al. [14], Ren et al. [16], Kirillov et al. [11] require the labels of the objects to perform instance segmentation. Image tagging models like Zhang et al. [18] can recognize the objects in a scene and return their labels, which can be used as text prompt to these models.
2.3 Representating the environment for task planning
@@
Prior works like Gu et al. [19], Chen et al. [20], Huang et al. [21], Chaplot et al. [22], Zhang et al. [23], Chen et al. [24] focused on building efficient representations of the environment to facilitate downstream task planning. The representations generated can be used for understanding the semantics of the scene, thus allowing the agent to query the environment for an object using natural language Gu et al. [19], Chen et al. [20] and generate task level plans Rana et al. [25], Chen et al. [20] or low-level actions Chen et al. [24]. Even though many scene representations like Chaplot et al. [22], Chen et al. [24] have created a scalable and efficient representation that can be used for navigation, it lacks contextual information about the scene, which is invaluable when exploring a scene. Scene representations like Gu et al. [19], Chen et al. [20] have a really dense representation, which is often not required when exploring an unfamiliar environment in search of a target object $G_{o}$. Our present research is inspired by the human way of cognition with an affordance based memory that attends and stores only those information which are relevant to the task at hand. We leverage on the ability of In-context learning in LLMs to prune out irrelevant information and then using an LVLM to reason about the relevant pruned list of objects, thus creating a sparse as well as information rich representation of the environment. The scene representation near the target object $G_{o}$ is made denser so as to facilitate further downstream task planning like manipulation of the object.
2.4 Object goal navigation
@@
Object goal navigation task involves finding a target object $G_{o}$ in an unfamiliar environment. Efficient exploration through the environment requires deep understanding of semantic priors of the world. Prior works like Chaplot et al. [22], Al-Halah et al. [26], Mousavian et al. [27], Chang et al. [28], Liang et al. [29], Wu et al. [30] tried to learn these semantic cues from egocentric RGB and depth images Al-Halah et al. [26], Mousavian et al. [27], semantic map Chaplot et al. [22] and even from YouTube videos Chang et al. [28]. These learned semantic cues might not generalize well to new unseen environments. Al-Halah et al. [26] focused on using a modular transfer learning model to generalize the learned policy for a particular task to multiple tasks and Wortsman et al. [31] used meta learning to generalize to unfamiliar environments. LLMs on the other hand exhibit extraordinary contextual awareness and ability to understand the semantic cues of our world. We leverage on this capability to efficiently explore unseen environments for the target object $G_{o}$.
@@
Recent works like Rajvanshi et al. [32], Dorbala et al. [33], Shah et al. [34] have leveraged the capability of LLMs to understand the semantic priors for object goal navigation tasks. Rajvanshi et al. [32] leveraged the planning capabilities of LLMs to devise a sequential plan that includes point goals and target state information for executing multiple object goal navigation tasks in parallel. These point goals generated by an LLM is fulfilled through a low level(execution level) controller. Shah et al. [34] focused on using the semantic predictions from LLMs as a heuristic for a frontier based exploration strategy to find the target object $G_{o}$. Dorbala et al. [33] focused on solving language-driven zero-shot object goal navigation problem Majumdar et al. [35] by using an LLM to navigate to the target object $G_{o}$, given a natural language description of it.
@@
3 METHOD
Figure 2: Architecture of the proposed pipeline.The agent explores the environment and collects observations (RGBD image and Pose). An open-vocabulary segmentation module is used to identify the objects in the current frame. Pruner takes in these detected segments and prunes out unwanted segments. The pruned segments are either initialized as a new node in the 3D scene representation or merged with an existing one, based on a similarity criteria. All new nodes are captioned with LLaVA, to provide semantic information to the LLM based planner. The agent then chooses a node to explore closer, to find the target object $G_{o}$. While doing so, the agent stores frame wise information in the short term memory module. If the agent decides that no objects in the 3D scene representation has a good chance of finding the target object $G_{o}$ closer to it, it continues to explore the scene and build the 3D scene representation.
@@ 3 METHOD
An overview of our approach is given in figure 2. The agent receives RGB image, depth image, base and camera pose as inputs in every timestep. An open vocabulary semantic segmentation module identifies objects in the RGB image and generates mask for the same. LLM acting as a pruner identifies the most important objects in the scene in regards to their utility in understanding the semantic priors of the environment and prunes the remaining detections. These detected segments are then converted into a 3D scene representation, where each object is signified by a node. An object node stores information regarding the position of the object, pointcloud, frame in which this object was detected, mask and label of this object in that frame, and a semantic description of the object in the scene. LLM, grounded for an object goal navigation task, uses the 3D scene modular representation to decide whether to continue exploring the environment or to move closer to an already detected object in the scene that has a high probability of finding the target object $G_{o}$. The agent triggers its short term memory module and stores information for every frame if it decides to explore closer to an object. Upon reaching the object, the agent utilize the detections from the short term memory, to construct hypotheses regarding the target object.
3.1 Open vocabulary image segmentation
@@
The open vocabulary segmentation module segments the RGB frame $I_{t}$ at the current timestep $t$, given the natural language description of the object. Given an RGB image $I_{t}$ of the scene, we use a combination of Recognize anything model(RAM) Zhang et al. [18], Grounding Dino and FastSAM Ren et al. [16] for open vocabulary semantic segmentation. RAM is a strong foundational model used for image tagging, it is tasked with finding the object tags $c_{t}^{i}$ in the scene. The set of object tags $C_{t}=(c_{t}^{1},c_{t}^{2},...,c_{t}^{m})$ is then passed as a text prompt $T_{t}$ $(=C_{t})$ to Grounding-Dino Liu et al. [17], which detects the presence of classes $C_{t}$ in the scene. Grounding-Dino returns the class labels $(l_{t}^{1},l_{t}^{2},...,l_{t}^{m})$ of objects in the scene, confidence of the predictions $(p_{t}^{1},p_{t}^{2},...,p_{t}^{m})$ and their bounding boxes $(b_{t}^{1},b_{t}^{2},...,b_{t}^{m})$. The detected classes in the scene $(l_{t}^{1},l_{t}^{2},...,l_{t}^{m})$ are then passed as box prompt $(b_{t}^{1},b_{t}^{2},...,b_{t}^{m})$ through FastSAM Zhao et al. [15] to get their corresponding masks $(m_{t}^{1},m_{t}^{2},...,m_{t}^{m})$.
3.2 LLM as a pruner
@@
We leverage the In-context learning abilities of an LLM to prune the detected class labels $(l_{t}^{1},l_{t}^{2},...,l_{t}^{m})$ generated by the open vocabulary segmentation module for the timestep $t$. In-context learning is a paradigm that empowers pre-trained LLMs to ground to a particular task without fine-tuning the model, by providing the LLM with task level demonstrations along with the prompt. The language model here is provided with task-specific examples containing a pair of input $P_{i}$ and output $P_{o}$ object sets. The input set $P_{i}$ contains class labels $(l^{1},l^{2},...,l^{m})$ and the output set $P_{o}(\subseteq P_{i})$, contains those objects that have a higher utility when it comes to understanding the semantic priors of the scene. The text prompt provided to the LLM to achieve this behaviour is given in fig LABEL:fig:pruner.
3.3 3D scene modular representation
@@
The pruned semantic masks $(m_{t}^{1},m_{t}^{2},...,m_{t}^{m})$ and labels $(l_{t}^{1},l_{t}^{2},...,l_{t}^{m})$ from the open vocabulary segmentation module, along with the depth image $I_{t}^{D}$ and camera pose $P_{t}$ in the current time step $t$ are used to generate 3D object nodes $(N_{c1},N_{c2},...,N_{ck})$ for each object $O_{i}$ in that frame $(I_{t},I_{t}^{D})$ using a modified concept graphs pipeline Gu et al. [19]. In concept graphs, the 3D object nodes $N_{ci}$ in the scene representation are fused from multiple views considering their spatial overlap, using an object association strategy that compares the semantic and visual similarity of these objects. Merging objects from different frames often creates conflicts in class labels, as the object segment can be associated with different class labels in different frames. We propose to use a Large Vision Language Model (LVLM) to resolve this conflict. The LVLM is tasked to find the class label that best suits the object detected, by giving it a cropped image of the detected object and the conflicting class labels.
@@
(a) Prompt to LLM for pruning the detections
(b) Prompt to LLM for planning.
3.4 LLM as a planner
LLM, grounded in the current task using task level description of the scene is used to generate a high level plan for the agent. Given the current objects in the environment (objects in the constructed 3D scene representation ($N_{c1},N_{c2},...,N_{ck}$)) and the generated descriptions for these object nodes, the LLM is tasked to identify whether to continue exploring the environment randomly and build up the 3D scene representation, or to explore the regions closer to an object node $N_{cj}$. If the LLM decides to explore the scene more, then $action$ is set to $<explore\textunderscore scene>$ and if it believes that there is a high chance of finding the target object $G_{o}$ near any of the detected objects, then $action$ is set to $<explore\textunderscore obj>$. This ability of an LLM to extract semantic relationships between the target object $G_{o}$ and objects in the scene ($N_{c1},N_{c2},...,N_{ck}$) helps the agent to explore efficiently in the environment. The text prompt provided to the LLM to achieve this behaviour is given in fig LABEL:fig:planner.
3.5 Short term memory(STM) and reasoning